# ValveFit: An analysis-suitable B-spline-based surface fitting framework for patient-specific modeling of tricuspid valves

### v1

@@ 3.2.2 The Loss Function
The tangent vectors, $\mathbf{T}_{u}$ and $\mathbf{T}_{v}$, and the surface unit normal vector, $\mathbf{N}$, are computed from the Jacobian matrix as follows:
$\mathbf{T}_{u}(u,v)$ $=\frac{\partial\mathbf{S}}{\partial u}=\begin{bmatrix}\dfrac{\partial x}{\partial u}\ \dfrac{\partial y}{\partial u}\ \dfrac{\partial z}{\partial u}\end{bmatrix}^{T}\,,$ (9)
$\mathbf{T}_{v}(u,v)$ $=\frac{\partial\mathbf{S}}{\partial v}=\begin{bmatrix}\dfrac{\partial x}{\partial v}\ \dfrac{\partial y}{\partial v}\ \dfrac{\partial z}{\partial v}\end{bmatrix}^{T}\,,$ (10)
$\mathbf{N}(u,v)$ $=\frac{\mathbf{T}_{u}(u,v)\times\mathbf{T}_{v}(u,v)}{\left\|\mathbf{T}_{u}(u,v)\times\mathbf{T}_{v}(u,v)\right\|}.$ (11)
To improve the local orthogonality between the $u$ and $v$ parametric directions in the physical domain and reduce excessive folding of the surface, we define the orthogonality regularization term, $R_{\text{orth}}(\mathbf{S})$, as
$R_{\text{orth}}(\mathbf{S})=\frac{1}{M}\sum_{k=1}^{M}\left\|\frac{\mathbf{t}_{k}^{u}\cdot\mathbf{t}_{k}^{v}}{\left\|\mathbf{t}_{k}^{u}\right\|\left\|\mathbf{t}_{k}^{v}\right\|}\right\|,$ (12)
where $\mathbf{t}_{k}^{u}=\mathbf{T}_{u}(u_{k},v_{k})$ and $\mathbf{t}_{k}^{v}=\mathbf{T}_{v}(u_{k},v_{k})$ denote the tangent vectors evaluated at each sampled parametric point $(u_{k},v_{k})$, and $M$ is the number of sampled points in the parametric domain. By calculating the cosine of the angle between the tangent vectors, $R_{\text{orth}}$ avoids excessive skewing or twisting of surface elements during deformation. By controlling the variation of the vertical ($z$) component of the unit normal vector $\mathbf{N}(u,v)$, the second term in Eq. 7, $R_{\text{norm}}(\mathbf{S})$, penalizes self-intersections of the template surface and is defined as
$R_{\text{norm}}(\mathbf{S})=\max_{1\leq k\leq M}\left\|\mathbf{n}_{k}^{z}-\bar{\mathbf{n}}^{z}\right\|,$ (13)
where $\mathbf{n}_{k}^{z}$ denotes the $z$-component of the unit normal vector $\mathbf{N}(u,v)$ evaluated at the parametric coordinate $(u_{k},v_{k})$, and $\bar{\mathbf{n}}^{z}$ is the mean value of the $z$-component of the normal vector over all the sampled parametric points. In order to prevent abrupt changes in surface orientation and intersections in the fitted surfaces, this term penalizes significant deviations in surface normal orientation along the ($z$) direction. Fig. 5 illustrates how the regularization terms in $\mathcal{E}_{\text{reg}}$ affect surface fitting results. In Fig. 5(b) the surface is fitted without the orthogonality term $R_{\text{orth}}$, resulting in severe folding of the surface— although no self-intersections occur. In contrast, fitting the surface without $R_{\text{norm}}$ leads to self-intersections, as seen in Fig. 5(c). When both $R_{\text{orth}}$ and $R_{\text{norm}}$ are included in the loss function, the resulting surface is free from severe folding and self-intersections (as shown in Fig. 5(a)).
Figure 5: B-spline surfaces fitted: (a) with both $R_{\text{orth}}$ and $R_{\text{norm}}$ included in the loss function, resulting in a smooth surface free from severe folding or self-intersections; (b) without $R_{\text{orth}}$, leading to severe folding due to loss of orthogonality of parametric lines; and (c) without $R_{\text{norm}}$, resulting in visible self-intersections due to large variations in surface normal orientation.
3.2.3 Surface Fitting
To accurately fit the template B-spline surface to target point clouds obtained from image segmentation, we use differentiable programming [64, 65] during surface fitting. In the template B-spline surface (Eq. (1)), the parametric coordinates $u$ and $v$ are fixed during optimization, and the control point positions $\mathbf{P}$ are optimized to minimize the loss function. The gradient of the loss function with respect to $\mathbf{P}$ is given as
@@ 3.2.3 Surface Fitting
where $\mathbf{s}_{k}$ are points sampled from the B-spline surface and $\mathbf{q}_{l}$ are points from the target point cloud. Manual gradient evaluation is laborious and ineffective since it would be necessary to derive and implement the gradients for each term in the loss function, which would involve significant bookkeeping and be prone to error. This complexity emphasizes how crucial it is to use automatic differentiation in our surface fitting framework in order to compute gradients effectively and precisely. The optimization of control points $\mathbf{P}$ to minimize the loss function is performed using the Adam optimizer [75]. The gradient of the loss function with respect to all control points is given by
@@
where $\delta$ is the learning rate and $t$ denotes the time step. The template B-spline surface $\mathbf{S}_{0}$ is initialized with predefined knot vectors and polynomial degrees at the start of the optimization process. The target point clouds, representing valve configurations across the cardiac cycle, are sequentially ordered from the open to the closed state, or vice versa. To ensure consistency in scale, position, and orientation, the template surface is first globally aligned to the corresponding target point clouds using affine transformations. In order to accurately fit $\mathbf{S}_{0}$ to $\mathbf{Q}_{1}$, non-linear deformation of the template surface is carried out by minimization of the loss function (Eq. (2)). The resulting fitted surface, $\mathbf{S}_{1}$, serves as the updated template surface for fitting the next point cloud in the sequence, $\mathbf{Q}_{2}$. This process allows for consistent surface fitting results throughout the cardiac cycle and is carried out throughout the entire sequence. The detailed algorithm is outlined in Algorithm 1.
@@
Input: Template geometry with predefined parametrization: basis functions $(\mathbf{B}_{u},\mathbf{B}_{v})$ and control points $\mathbf{P}_{0}$; sequence of target point clouds $\{\mathbf{Q}_{1},\mathbf{Q}_{2},\dots,\mathbf{Q}_{n}\}$; maximum number of iterations $t_{\text{max}}$; learning rate $\delta$; weights $\{w_{\text{cd}},w_{\text{hd}},w_{\text{a}},w_{\text{orth}},w_{\text{norm}}\}$
@@ 4 Results
To quantitatively evaluate the surface fitting accuracy, we introduce a dimensionless error metric, the scaled Nearest Neighbor Distance (sNND). While the Nearest Neighbor Distance (NND) measures the minimum distance from each point in the point cloud to the fitted surface, it is scale-dependent, varying with the size of the geometry. To address this limitation, sNND normalizes this nearest neighbor distance by the square root of the surface area, allowing for consistent comparison across valve geometries of different anatomical sizes in patients. For each point $\mathbf{q}_{l}$ in the target point cloud, where $l=1,\cdots,N$, the sNND to the fitted surface $\mathbf{S}$ of area $\mathcal{A}(\mathbf{S})$ is defined as
@@ 4.1 Validation Study with Simulation Data
These structural simulations are carried out using an IGA-based tricuspid valve parameterization framework [20]. This framework defines the annulus during the systolic phase of the cardiac cycle and determines the commissure and leaflet locations along the annular curve. These locations are then used to define the valve heights, whose endpoints are interpolated to construct the free-edge curve of the leaflet. From the annular and free-edge curves, a Gordon surface [71] is generated to reconstruct the leaflets’ surface parameterization. For the chordae tendineae, we define the positions of the papillary muscles in $R^{3}$ space, by specifying the number and spacing of the chordae. A steady-state structural simulation is performed on a dense analysis mesh by applying a follower pressure load normal to the leaflet surface, continuing until valve closure. Further details on the surface parameterization and simulation methodology can be found in [20]. The resulting simulation-derived point clouds represent idealized yet exact, noise-free surfaces, where accurate fitting results at different stages of valve motion are expected. To match the point cloud density typically observed in image segmentations of patient data, we sample approximately $1500$ points from the simulation surface for fitting. This ensures that the benchmark point cloud data has a similar point cloud density as the patient data.
@@
Figure 6: The fitted B-spline surfaces are shown overlaid with the corresponding benchmark point clouds, obtained at multiple steps during the structural simulation of valve motion: (a) Step 0, (b) Step 30, (c) Step 40, (d) Step 50, and (e) Step 60. The point clouds are colored based on the sNND error with respect to the fitted surface.
@@
The template geometry shown in Fig. 2 is first re-parameterized using uniform knot vectors, resulting in 23 control points with degree $p=3$ along the $u$ direction and 5 control points with degree $q=2$ along the $v$ direction. The template surface is then aligned to the target point cloud at simulation step 00 (Fig. 6(a))using affine transformations. The positions of the control points are then updated by surface fitting, where each fitted B-spline surface acts as a template for the next point cloud in the sequence. The weight parameters in the loss function (Eq. (2)) were set to $w_{\text{cd}}=20$, $w_{\text{hd}}=0.5$, $w_{\text{a}}=0$, $w_{\text{orth}}=10$, and $w_{\text{norm}}=5$. These weights are constant during the sequential fitting of the B-spline surface. The weight $w_{\text{a}}$ is set to zero because the point cloud for the validation study is not further segmented into annulus and leaflet regions. Therefore, this dataset does not have explicit labels for the points belonging to the annulus and leaflets. In Fig. 6, we show the fitted B-spline surfaces overlaid with the corresponding point cloud data at each simulation step. The sNND error is computed for each point in the point cloud to evaluate local fitting accuracy. Based on the results, we can see that the resulting surfaces are smooth, intersection-free, and closely aligned with the point clouds across all the stages of valve motion. Complex folding structures of the geometry are effectively captured, while maintaining smoothness and surface regularity.
@@
Figure 7: Histograms showing the percentage of points with respect to sNND error at different simulation steps for the validation study. The distributions are consistent across all steps, with the majority of points exhibiting low fitting error, concentrated in the range below $0.02$. This indicates the stable performance of the ValveFit framework over time.
@@
00 $3.06\times 10^{-4}$ $0.0363$ $0.0087$
$30$ $2.23\times 10^{-4}$ $0.0430$ $0.0091$
$40$ $5.66\times 10^{-5}$ $0.0353$ $0.0087$
$50$ $1.07\times 10^{-4}$ $0.0343$ $0.0103$
$60$ $6.38\times 10^{-5}$ $0.0342$ $0.0109$
The histograms shown in Fig. 7 show consistent sNND error distributions across all the simulation steps, with the majority of points concentrated in the low sNND range (below $0.02$). Table 1 presents a quantitative assessment of the surface fitting accuracy using the sNND error metric, where lower values indicate better alignment between fitted surfaces and point cloud data. Here, we report the minimum, mean, and maximum sNND errors for each fitting result. The mean sNND error quantifies the overall fitting accuracy, while the difference between maximum and minimum sNND errors quantifies the amount of variability in the fitting accuracy. Across all the simulation steps, low sNND error values are observed, with minimum sNND values ranging between $5.66\times 10^{-5}$ and $3.06\times 10^{-4}$. In all the simulation steps, the maximum sNND values stay below $0.0430$. The mean sNND values range between $0.0087$ and $0.0109$, showing consistent fitting at all stages of valve motion. We can see that high accuracy is maintained across most of the geometry, with maximum sNND values localized to regions of greater complexity, particularly near leaflet edges and areas undergoing folding. The framework effectively balances geometric accuracy with the requirement of maintaining intersection-free surfaces throughout the cardiac cycle.
@@ 4.2 Surface fitting sensitivity with respect to point cloud sampling
Figure 8: Fitted B-spline surfaces overlaid with target point clouds obtained at simulation step $60$ (see Fig. 6(e)), sampled at varying densities ((a)(f): $500$ to $5000$ points). Higher point cloud densities provide more detailed geometric information, enabling more accurate fitting results.
Based on the results shown in Table 2 and Fig. 8, we observe that the sNND error continuously decreases from $0.0155$ to $0.0121$ as the point cloud density increases from $50$ to $5000$. This is because, in order to capture the intricate details of the valve geometry during fitting, a higher point cloud density is required to provide a more detailed geometric representation. However, we observe that the mean sNND error does not decrease further beyond a point cloud density of $2500$ points. This is because the fitted B-spline surface is evaluated at $2500$ points for the calculation of the sNND metric. When calculating the Chamfer distance—which involves nearest neighbor distances from the point cloud to the surface points and vice versa—the resolution of the surface sampling becomes the limiting factor. Further improvements in the sNND metric due to an increase in sampling have no noticeable impact on nearest neighbor distance values once the point cloud density becomes greater than the number of points sampled on the B-spline surface. The percentage of points in the lower error bracket increases with point cloud density, as can be seen in Fig. 9 and Table 2. As point cloud density increases, the difference between the maximum and minimum sNND errors also decreases. This is because a higher point cloud density reduces the chance of overfitting by incorporating more point cloud-to-surface distance evaluations into the Chamfer distance.
@@
$50$ $5.06\times 10^{-5}$ $0.0718$ $0.0155$
$250$ $5.70\times 10^{-5}$ $0.0528$ $0.0132$
$500$ $8.03\times 10^{-5}$ $0.0396$ $0.0124$
$1000$ $1.40\times 10^{-4}$ $0.0382$ $0.0121$
$2500$ $8.85\times 10^{-5}$ $0.0364$ $0.0119$
$5000$ $4.02\times 10^{-4}$ $0.0374$ $0.0121$

### v2

@@ 4.2.2 Surface Regularization Energy
The tangent vectors ($\mathbf{t}^{u}$ and $\mathbf{t}^{v}$) and the unit surface normal vector $\mathbf{n}$ are derived from the Jacobian matrix as follows:
$\mathbf{t}^{u}(u,v)$ $=\frac{\partial\mathbf{S}}{\partial u}=\begin{bmatrix}\dfrac{\partial x}{\partial u}\ \dfrac{\partial y}{\partial u}\ \dfrac{\partial z}{\partial u}\end{bmatrix}^{T}\,$, (12)
$\mathbf{t}^{v}(u,v)$ $=\frac{\partial\mathbf{S}}{\partial v}=\begin{bmatrix}\dfrac{\partial x}{\partial v}\ \dfrac{\partial y}{\partial v}\ \dfrac{\partial z}{\partial v}\end{bmatrix}^{T}\,$, (13)
$\mathbf{n}(u,v)$ $=\frac{\mathbf{t}^{u}(u,v)\times\mathbf{t}^{v}(u,v)}{\left\|\mathbf{t}^{u}(u,v)\times\mathbf{t}^{v}(u,v)\right\|}.$ (14)
4.3 Surface Fitting
To accurately fit the template B-spline surface to target point clouds obtained from image segmentation, we use differentiable programming during surface fitting [73, 74]. In the template B-spline surface (Eq. (1)), the parametric coordinates $u$ and $v$ are fixed during optimization, and the control point positions $\mathbf{P}$ are optimized to minimize the loss function. The gradient of the loss function with respect to $\mathbf{P}$ is given as
@@ 4.3 Surface Fitting
where $\mathbf{s}_{k}$ are points sampled from the B-spline surface and $\mathbf{q}_{l}$ are points from the target point cloud. Manual gradient evaluation is laborious and ineffective since it would be necessary to derive and implement the gradients for each term in the loss function, which would involve significant bookkeeping and be prone to error. This complexity emphasizes how crucial it is to use automatic differentiation in our surface fitting framework in order to compute gradients effectively and precisely. The gradient of the loss function with respect to all control points is given by
@@
where $\delta$ is the step-size and $t$ denotes the time step. We have utilized a step-size of $0.001$, determined empirically based on the maximum number of iterations and rate of convergence. The optimization of control points $\mathbf{P}$ to minimize the loss function is performed using the Adam optimizer [87].
The surface fitting process begins by initializing a template B-spline surface ($\mathbf{S}_{0}$) with predefined basis functions $(\mathbf{B}_{u},\mathbf{B}_{v})$ and initial control points ($\mathbf{P}_{0}$) at the start of the optimization process. The target data consist of a sequence of point clouds $\{\mathbf{Q}_{1},\mathbf{Q}_{2},\dots,\mathbf{Q}_{n}\}$, each segmented from echocardiographic images at various stages of cardiac cycle. These point clouds are ordered from the fully open state to the fully closed state.
For each point cloud in the sequence, the fitting process starts by initializing the control points of the surface with those from the previous stage. For the first point cloud $\mathbf{Q}_{1}$, the process begins with the control points of initial template $\mathbf{P}_{0}$. To provide a good initial alignment, an affine transformation is applied to the surface so that it matches the global position, scale, and orientation of the target point cloud. The template surface is iteratively deformed to match the target point cloud driven by the gradient of the loss function (Eq. (19)). A step size $\delta=0.001$ and a maximum of $t_{max}=10,000$ iterations were employed to ensure stable convergence and accurate surface fitting.
Upon convergence for a given point cloud, the optimized control points are stored. These control points are then used to initialize the fitting process for the subsequent point cloud. Repeating this procedure across the entire sequence yields a series of control points $\{\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{n}\}$ corresponding to the fitted surfaces $\{\mathbf{S}_{1},\mathbf{S}_{2},\dots,\mathbf{S}_{n}\}$. The program outlined in Algorithm 1 is implemented utilizing JAX arrays, which runs array operations on both GPU and CPU.
@@
Input: Template geometry with predefined parametrization: basis functions $(\mathbf{B}_{u},\mathbf{B}_{v})$ and control points $\mathbf{P}_{0}$; sequence of target point clouds $\{\mathbf{Q}_{1},\mathbf{Q}_{2},\dots,\mathbf{Q}_{n}\}$; maximum number of iterations $t_{\text{max}}$; step-size $\delta$; weights $\{w_{\text{CD}},w_{\text{HD}},w_{\text{a}},w_{\text{orth}},w_{\text{norm}}\}$
@@ 5 Results
To quantitatively evaluate the surface fitting accuracy, we introduce a dimensionless error metric, the scaled Nearest Neighbor Distance (sNND). While the Nearest Neighbor Distance (NND) measures the minimum distance from each point in the point cloud to the fitted surface, it is scale-dependent, varying with the size of the geometry. To address this limitation, sNND normalizes this nearest neighbor distance by the square root of the surface area, allowing for consistent comparison across valve geometries of different anatomical sizes in patients. For each point $\mathbf{q}_{l}$ in the target point cloud, where $l=1,\cdots,N$, the sNND to the fitted surface of area $\mathcal{A}(\mathbf{S})$ is defined as
@@
The following results were generated on a hardware configuration consisting of an $\text{Intel}^{\text{\textregistered}}$ $\text{Core}^{\text{TM}}$ i9-14900KF CPU, 64 GB RAM, and an $\text{Nvidia}^{\text{\textregistered}}$ GeForce $\text{RTX}^{\text{TM}}$ 4090 GPU.
@@ 5.1 Validation Study with Simulation Data
These structural simulations are carried out using an IGA-based tricuspid valve parameterization framework [20]. This framework defines the annulus during the systolic phase of the cardiac cycle and determines the commissure and leaflet locations along the annular curve. These locations are then used to define the valve heights, whose endpoints are interpolated to construct the free-edge curve of the leaflet. From the annular and free-edge curves, a Gordon surface [80] is generated to reconstruct the leaflets’ surface parameterization. For the chordae tendineae, we define the positions of the papillary muscles in $R^{3}$ space, by specifying the number and spacing of the chordae. A steady-state structural simulation is performed on a dense analysis mesh by applying a follower pressure load normal to the leaflet surface, continuing until valve closure. Further details on the surface parameterization and simulation methodology can be found in [20]. The resulting simulation-derived point clouds represent idealized yet exact, noise-free surfaces, where accurate fitting results at different stages of valve motion are expected. To match the point cloud density typically observed in image segmentations of patient data, we sample approximately $1500$ points from the simulation surface for fitting. The sampling is performed using the point cloud simplification filter in MeshLab based on the Poisson disk strategy [88, 89].
@@
Figure 11: The fitted B-spline surfaces are shown overlaid with the corresponding benchmark point clouds, obtained at multiple steps during the structural simulation of valve motion: (a) Step 30, (b) Step 50, (c) Step 70, and (d) Step 90. The point clouds are colored based on the sNND error with respect to the fitted surface.
@@
The template geometry shown in Fig. 2 is first re-parameterized using uniform knot vectors, resulting in $33$ control points with degree $p=3$ along the $u$ direction and $6$ control points with degree $q=3$ along the $v$ direction. The template surface is then aligned to the target point cloud at simulation step $30$ (Fig. 11(a)) using affine transformations. The positions of the control points are then updated by minimizing the loss function, and the resulting fitted surface is used as the template surface for the next point cloud in the cardiac cycle. The weight parameters in the loss function (Eq. (2)) were set to $w_{\text{CD}}=80$, $w_{\text{HD}}=30$, $w_{\text{a}}=0$, $w_{\text{orth}}=5$, $w_{\text{TPE}}=10^{-7}$, and $w_{\text{norm}}=2$. These weights are constant during the sequential fitting of the target point clouds. The value of $w_{\text{TPE}}$ is very low because the TPE term is inherently scale-dependent. In the patient dataset (Section 5.4), the geometries are larger in scale than those in the validation dataset (Section 5.1), resulting in numerically larger inter-point distances. Since TPE values decay sharply with increasing distance (Fig. 9(b)), we increase the $w_{\text{TPE}}$ value to maintain an effective repulsive force for larger inter-point distances. The weight $w_{\text{a}}$ is set to zero because this dataset does not have explicit labels for the points belonging to the annulus and leaflets. We utilize the same weights for the results presented in the Sections 5.25.3 as well.
In Fig. 11, we show the fitted B-spline surfaces overlaid with the corresponding point cloud data at each simulation step. The sNND error is computed for each point in the point cloud to evaluate local fitting accuracy. Based on the results, we can see that the fitted surfaces are smooth, intersection-free, and closely aligned with the point clouds across all the stages of valve motion. Complex folding structures of the geometry are effectively captured, while maintaining smoothness and surface regularity (Fig. 11(c)-(d)).
@@
Figure 12: Histograms showing the percentage of points with respect to sNND error at different simulation steps for the validation study. The distributions are consistent across all steps, with the majority of points exhibiting low fitting error, concentrated in the range below $0.005$. This indicates the stable performance of the ValveFit framework over time.
@@
$30$ $8.25\times 10^{-5}$ $6.43\times 10^{-3}$ $2.51\times 10^{-3}$
$50$ $3.87\times 10^{-5}$ $7.66\times 10^{-3}$ $3.47\times 10^{-3}$
$70$ $1.06\times 10^{-4}$ $9.21\times 10^{-3}$ $4.35\times 10^{-3}$
$90$ $4.37\times 10^{-5}$ $1.19\times 10^{-2}$ $5.38\times 10^{-3}$
The histograms shown in Fig. 12 show consistent sNND error distributions across all the simulation steps, with the majority of points concentrated in the low sNND range (below $0.005$). Table 1 presents a quantitative assessment of the surface fitting accuracy using the sNND error metric, where lower values indicate better alignment between fitted surfaces and the target point cloud. Here, we report the minimum, mean, and maximum sNND errors for each fitting result. The mean sNND error quantifies the overall fitting accuracy, while the difference between the maximum and minimum sNND errors quantifies the amount of variability in the fitting accuracy. Across all the simulation steps, low sNND error values are observed, with minimum sNND values ranging between $3.87\times 10^{-5}$ and $1.06\times 10^{-4}$. In all the simulation steps, the maximum sNND values stay below $0.0119$. The mean sNND values range between $2.51\times 10^{-3}$ and $5.38\times 10^{-3}$, showing consistent fitting at all stages of valve motion. We can see that high accuracy is maintained across most of the geometry, with maximum sNND values localized to regions of greater complexity, particularly near leaflet edges and areas undergoing coaptation as seen in Fig. 11(d). The fitted surface also captures the narrow folds seen in Fig. 11(c)-(d). The framework effectively balances geometric accuracy with the requirement of maintaining intersection-free surfaces throughout the cardiac cycle.
@@ 5.2 Surface Fitting Sensitivity with Respect to Point Cloud Sampling
Figure 13: Fitted B-spline surfaces overlaid with target point clouds obtained at simulation step $60$, sampled at varying densities ((a)(f): $50$ to $5000$ points).
@@
Based on the results shown in Table 2 and Fig. 14, we observe that the mean sNND error increases from $6.45\times 10^{-5}$ to $0.00376$ as the point cloud density increases from 50 to 2500 points. At a low density of 50 points, the sparse sampling inadequately captures the complexity of the valve geometry, leading to an artificially low sNND error during fitting. As the point cloud density increases, it better represents the intricate geometric features of the valve. Thus, the surface fitting framework needs to resolve more complex geometric features, resulting in a higher and more representative mean error. We note that when calculating the one-sided Chamfer distance—which involves nearest neighbor distances from the point cloud to the surface points—the density of template surface points becomes the limiting factor. Further improvements in the sNND metric due to an increase in point cloud density have no noticeable impact on nearest neighbor distance values once the point cloud density becomes greater than the number of points sampled on the B-spline surface. The results show that the overall topology of the fitted B-spline surface is established even at low point cloud densities. While finer geometric features are resolved as point density increases, the underlying shape of the surface fitted to just 50 points remains qualitatively consistent with surfaces obtained from denser point clouds.
@@
$50$ $2.55\times 10^{-6}$ $1.92\times 10^{-3}$ $6.45\times 10^{-5}$
$250$ $5.70\times 10^{-6}$ $1.46\times 10^{-2}$ $3.15\times 10^{-3}$
$500$ $2.01\times 10^{-5}$ $1.05\times 10^{-2}$ $3.50\times 10^{-3}$
$1000$ $3.07\times 10^{-5}$ $8.75\times 10^{-3}$ $3.57\times 10^{-3}$
$2500$ $4.11\times 10^{-5}$ $8.43\times 10^{-3}$ $3.76\times 10^{-3}$
$5000$ $7.03\times 10^{-5}$ $8.35\times 10^{-3}$ $3.82\times 10^{-3}$